OpenStudy (anonymous):

PERCENTAGES QUESTION :) Two beakers of equal volume Beaker A contains 30% salt Beaker B contains 10% salt How many ounces of solution A should be added to solution B, so that the resulting solution is 20% salt?

OpenStudy (anonymous):

so far, I let x be the amount of solution in Beaker A. then, 30% of that is the amount of salt in A....same for B...

OpenStudy (mathstudent55):

Let's try 10 liters. 10 liters of A has 3 litres salt. 10 liters of B has 1 litres salt. You end up with 20 liters with 4 litres salt. 4 litres out of 20 litres is 20%

OpenStudy (anonymous):

Okay, yep, I;m following so far....

OpenStudy (mathstudent55):

Since the concentration you want is exactly the average of the two concentrations you start with, the same amounts of A and B are used.

OpenStudy (anonymous):

Oki doki, that makes sense - there might be a problem with the worked solution – its talking about ounces, and I have no information about how many ounced the beakers started out with....

OpenStudy (mathstudent55):

Use ounces where I used litres.

OpenStudy (anonymous):

heres the part that dosnt make any sense to me in the worked solution: "the resulting solution will be 50 + x ounces and its concentration of salt will be 20%(50+x)....

OpenStudy (anonymous):

im guessing instead of your 10L base, theyve used 50 ounces?

OpenStudy (mathstudent55):

Is there any info missing from the question you posted? Where does the number 50 come from in the solution?

OpenStudy (anonymous):

thats exactly my problem...

OpenStudy (mathstudent55):

Beaker A has x ounces of 30% salt solution. Beaker B has x ounces of 10% salt solution. (Both beakers have x amount because we are the beakers have equal volume.)

OpenStudy (anonymous):

Yes. Corecct

OpenStudy (mathstudent55):

x ounces of solution B has 0.1x ounces of salt.

OpenStudy (anonymous):

Yep

OpenStudy (mathstudent55):

Now we need to start adding solution A to solution B until we have a new solution that is 20% salt.

OpenStudy (mathstudent55):

Let's say we add y ounces of solution A to solution B.

OpenStudy (anonymous):

To do that, we need to find the volume in A containing a concentration of 10%?

OpenStudy (anonymous):

Okay, yep

OpenStudy (mathstudent55):

y ounces of solution A has 0.3y ounces of salt

OpenStudy (anonymous):

correct

OpenStudy (mathstudent55):

We will have a total of salt: 0.3y + 0.1x Ok?

OpenStudy (anonymous):

Yep, all good so far. Thanks for going through it with me.

OpenStudy (mathstudent55):

Next, we will have a total amount of solution of: x + y Ok?

OpenStudy (anonymous):

Yes.

OpenStudy (mathstudent55):

We want the new solution to have a concentration of 20% salt. 20% = 0.2 That means \(\dfrac{0.3y + 0.1x}{x + y } = 0.2\)

OpenStudy (anonymous):

Okay, cool.

OpenStudy (mathstudent55):

Now let's solve that equation and what we get.

OpenStudy (anonymous):

x=y?

OpenStudy (mathstudent55):

\(\dfrac{0.3y + 0.1x}{x + y } = 0.2\) \(0.3y + 0.1x = 0.2(x + y)\) \(0.3y + 0.1x = 0.2x + 0.2y\) \(0.1y = 0.1x\) \(y = x\) Correct.

OpenStudy (mathstudent55):

Since we chose x to be the amount of solution B and y to be the amount of solution A, we see the amounts must be the same no matter how much each beaker has.

OpenStudy (anonymous):

Understood

OpenStudy (mathstudent55):

Great.

OpenStudy (anonymous):

So do you think that they made x=y=50 ounces?

OpenStudy (anonymous):

in the solutions?

OpenStudy (mathstudent55):

That's what I am thinking.

OpenStudy (anonymous):

Yeah, okay. Well thanks for going through it with me in detail. Appreciate your work :)

OpenStudy (mathstudent55):

Before the part of the solution you quoted above mentioning 50 + x, was there any mention of 50?

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

their equation is exactly what you had, except they made 50=y for some reason: 10%50 + 30%x = 20%(50+x) I understand the equation now, thank you. And no. No mention anywhere of 50 :(

OpenStudy (anonymous):

Also, just wondering if you know how to change the username on open study? i can find the change passwords under settings, but is it even possible to change usernames?

OpenStudy (mathstudent55):

I guess they just assumed they would use 50 ounces of one solution, and then they found how many ounces of the other solution they needed. It turns out to be 50, of course, bec now we know it's the same amount of both solutions.

OpenStudy (mathstudent55):

I don't know if you can change username in OS.

OpenStudy (anonymous):

Yes, I reckon you're right on that one. Okay then, thank you!